OpenStudy (jessiegonzales):

MEDAL!!! ALGEBRA 2 4. What is the slope of the line that passes throughout the given points? (3,2) and (5,12) 5 -5 -1/5 1/5

pooja195 (pooja195):

\[\huge~\rm~\frac{ y_2-y_1 }{ x_2-x_1 }=slope\]

OpenStudy (jessiegonzales):

2-12 3-5

pooja195 (pooja195):

No

OpenStudy (misty1212):

lol yes!

pooja195 (pooja195):

@misty1212 12-2

OpenStudy (misty1212):

\[\huge \frac{2-12}{3-5}\] for sure

jimthompson5910 (jim_thompson5910):

@pooja195 as long as you keep the order the same, it doesn't matter

OpenStudy (jessiegonzales):

-10/2 -5

pooja195 (pooja195):

3-5 is not 2

OpenStudy (misty1212):

god does not tell you which one is \((x_1,y_1)\) and \((x_2,y_2)\) you get to pick that for yourself

pooja195 (pooja195):

@misty1212 ok i get it.

OpenStudy (jessiegonzales):

my answer is -5

pooja195 (pooja195):

thats not right -10/-2=?

pooja195 (pooja195):

3-5=-2 NOT 2

jimthompson5910 (jim_thompson5910):

(x1,y1) = (3,2) (x2,y2) = (5,12) ------------------------------------------------------- \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{12 - 2}{5-3}\] \[\Large m = \frac{10}{2}\] \[\Large m = 5\] ------------------------------------------------------- \[\Large m = \frac{y_{1}-y_{2}}{x_{1}-x_{2}}\] \[\Large m = \frac{2-12}{3-5}\] \[\Large m = \frac{-10}{-2}\] \[\Large m = 5\] ------------------------------------------------------- Either way, you get the same slope.